if tan(A-B)=1, sec(A+B)=2/sq. root 3, then find the smallest positive value of B?
 

Grade:11

1 Answers

Arun
6 years ago
a-b=nPi+ Pi/41/cos(a+b)=sec(a+b)=2/root 3cos(a+b)=root 3/2a+b=2nPi+Pi/6a=2nPi+1/2(Pi/4+Pi/6)b=2nPi+1/2(Pi/6-Pi/4)for n=0,smallest value of a=1/2(Pi/4+Pi/6)smallest value of b=1/2(Pi/6-Pi/4)